OpenStudy (anonymous):

Find: (f(x+c)-f(x))/c If f(x) = 2x^3-5x

OpenStudy (freckles):

can you evaluate f(x+c) if given f(x)=2x^3-5x hint: just replace the old input x wherever you see with the new input (x+c)

OpenStudy (anonymous):

@freckles how would you incorporate c if we aren't given a value?

OpenStudy (freckles):

example: if f(x)=x^2 then f(x+c)=(x+c)^2

OpenStudy (freckles):

wherever you see x you replace it with (x+c)

OpenStudy (freckles):

if f(x)=5-x+x^2 then f(x+c)=5-(x+c)+(x+c)^2

OpenStudy (freckles):

so if f(x)=2x^3-5x then f(x+c)=?

OpenStudy (freckles):

once you find f(c+h) we can plug into difference quotient when you are ready for that

OpenStudy (anonymous):

@freckles 2(x+c)^3-5(x+c) ?

OpenStudy (freckles):

great! \[\frac{f(x+c)-f(c)}{h}=\frac{[2(x+c)^3-5(x+c)]-[2c^3-5c]}{h}\] now this is the part where we multiply everything out on the top and see if anything cancels in the top

OpenStudy (freckles):

oops that h down there is suppose to be a c by the way

OpenStudy (freckles):

and i read your problem wrong f(c) is f(x) lol

OpenStudy (anonymous):

Okay that's what I thought lol!

OpenStudy (freckles):

\[\frac{f(x+c)-f(x)}{c}=\frac{2(x+c)^3-5(x+c)]-[2x^3-5x]}{c}\]

OpenStudy (freckles):

now we still need to multiply stuff out on the top and see if anything cancels on the top

OpenStudy (freckles):

if i was multipying out (x+c)^3 i would refer to my handy dandy pascal's triangle

OpenStudy (freckles):



OpenStudy (freckles):